OpenStudy (anonymous):

Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x). x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 3 4 5 6 Just imagine like a table it wouldn't copy and paste.

OpenStudy (elonasushchik):

I'm just going to solve that equation: 2.5x – 10.5 = 64(0.5x) 2.5x - 10.5 = 32x 2.5x - 32x = 10.5 -29.5x = 10.5 x = -10.5/29.5

Nnesha (nnesha):

there are x values so replace x by its value and solve for y f(x) and g(x) is same as y

Nnesha (nnesha):

for example first one is 2 so replace x by 2 and then solve for y

OpenStudy (anonymous):

O OK that makes more since.

Nnesha (nnesha):

okays so when x =2 y = what ?

OpenStudy (anonymous):

-5.5

OpenStudy (anonymous):

?? @Nnesha

OpenStudy (anonymous):

@DivineSolar

Nnesha (nnesha):

yes right f(2) = -5.5 now replace x by 2 in this equation g(x) = 64(.5x) solve for g(2)

OpenStudy (anonymous):

g(2)=16

Nnesha (nnesha):

`16 how did you get that ?

OpenStudy (anonymous):

I'm sorry the g(x) equation the x is suppose to be a exponent

Nnesha (nnesha):

ohh okay so yes that's right now repeat the pattern!

OpenStudy (anonymous):

OK thanks so much

Nnesha (nnesha):

my pleasure :)